OpenStudy (anonymous):

find the inverse of f(x)=3x+2

OpenStudy (anonymous):

x=3f(x)+2 x-2=3f(x) f(x)=(x-2)/3 Swap the x and f(x) in the original equation

OpenStudy (anonymous):

your function says multiply by 3 and then add two inverse will say subtract 2, divide by 3 \[f^{-1}(x)=\frac{x-2}{3}\] just like if you were asked to solve \[3x+2=11\] you would first subtract 2 to get \[3x=9\] then divide by 3 to get \[x=3\] meaning if you know the output of the function you can find the input